OpenStudy (anonymous):

Choose one of the factors of 500x^3 + 108y^18 Answers: 25x^2 + 15xy^6 + 9y^2 All of the above 6 5x + 3y^6

OpenStudy (aroub):

5x + 3y^6

OpenStudy (aroub):

oh and i suggest u to go to this site: http://www.wolframalpha.com/

OpenStudy (anonymous):

supposed to recognize this as "difference of two cubes" \[a^3-b^2=(a-b)(a^2+ab+b^2)\]

OpenStudy (anonymous):

or rather "sum of two cubes" \[(a^3+b^3=(a+b)(a^2-ab+b^2)\]

OpenStudy (anonymous):

first factor out a 4 to get \[4(125x^3+27y^{18})\] and now \[a=5x,b=3y^6\] use the formula get \[4(5x+3y^6)(25x^2-15xy^6+9y^{12})\]

OpenStudy (anonymous):

so it looks like answer 1 is possible but it is wrong because of the sign